OpenStudy (aaronandyson):

2 cot A = 3 Find the value of 5 sin A - 2 cos A/6 cos A + sin A @ganeshie8 @Hero @iambatman @Directrix @butterflydreamer

OpenStudy (aaronandyson):

\[\frac{ 5 \sin A - 2 \cos A}{ 6 \cos A + \sin A }\]

OpenStudy (aaronandyson):

@Nnesha

Directrix (directrix):

cot A = cos A / sin A If 2 cot A = 3, then cot A = 3/2 and cos A / sin A = 3/2

Directrix (directrix):

On cos A / sin A = 3/2 , cross multiply to get 3 sin A = 2 cos A

Directrix (directrix):

Look at just the numerator of the fraction (5 sin A - 2 cos A)/(6 cos A + sin A)

Directrix (directrix):

(5 sin A - 2 cos A). Keep in mind that 3 sin A = 2 cos A and make substitutions. If 3 sin A = 2 cos A, then the numerator 5 sin A - 2 cos A = 5 sin A - 3 sin A = 2 sin A. See if you agree with that. @AaronAndyson Note that we still have to work with the denominator.

OpenStudy (aaronandyson):

Yes. I understood this, now what?

Directrix (directrix):

Please stay with me in this thread while we are working.

OpenStudy (aaronandyson):

okay.

Directrix (directrix):

If 3 sin A = 2 cos A, then 6 cos A + sin A = 9 sin A + sin A = 10 sin A in the denominator.

Directrix (directrix):

From above the numerator is: 2 sin A. The denominator is: 10 sin A 2 sin A / 10 sin A = 2/10 = 1/5

Directrix (directrix):

Look at that and see what you think.

OpenStudy (aaronandyson):

#justgotlost

OpenStudy (aaronandyson):

can you explain it again?

Directrix (directrix):

I have written out all of the steps. When I was doing most of them, I did not realize that you were observing another thread and were not here. I think that what I have written is clear enough for you to go back and read. If you have a specific question, then I will try to answer that.

OpenStudy (aaronandyson):

I was here all the time. The thing is I had a sleepless night.

Directrix (directrix):

I saw this which gave the idea that you were elsewhere.

OpenStudy (aaronandyson):

My lil bro must be using it :)

OpenStudy (aaronandyson):

and I keep multiple tabs of OS open.

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (aaronandyson):

@Nnesha

OpenStudy (aaronandyson):

@perl @Preetha

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (aaronandyson):

@Hero

OpenStudy (aaronandyson):

@ParthKohli